# DINO-WM: World Models on Pre-trained Visual Features enable Zero-shot Planning

### v1

@@ 3.1.3 Decoder for Interpretability
To aid in visualization and interpretability, we use a stack of transposed convolution layers to decode the patch representations back to image pixels, similar as in Razavi et al. (2019). Given a pre-collected dataset, we optimize the parameters $\theta$ of the decoder $q_{\theta}$ with a simple reconstruction loss defined as:
@@
The training of the decoder is entirely independent of the transition model training, offering several advantages: 1) The quality of the decoder does not affect the world model’s reasoning and planning capabilities for solving downstream tasks, and 2) During planning, there is no need to reconstruct raw pixel images, thereby reducing computational costs. Nevertheless, the decoder remains valuable as it enhances the interpretability of the world model’s predictions.
@@ 3.2 Visual Planning with DINO-WM
Arguably, to evaluate the quality of the world model, it needs to be able to allow for downstream reasoning and planning. A standard evaluation metric is to perform trajectory optimization at test time with these world models and measure the performance. While the planning methods themselves are fairly standard, it serves as a means to emphasize the quality of the world models. For this purpose, our world model receives the current observation $o_{0}$ and a goal observation $o_{g}$, both represented as RGB images. We formulate planning as the process of searching for a sequence of actions that the agent would take to reach $o_{g}$. To achieve this, we employ model predictive control (MPC), which facilitates planning by considering the outcomes of future actions.
We utilize the cross-entropy method (CEM), a stochastic optimization algorithm, to optimize the sequence of actions at each iteration. The planning cost is defined as the mean squared error (MSE) between the current latent state and the goal’s latent state, given by
$\mathcal{C}=\left\|\hat{z}_{T}-z_{g}\right\|^{2},\quad\text{where}\quad\hat{z}_{t}=p(\hat{z}_{t-1},a_{t-1}),\quad\hat{z}_{0}=\text{enc}(o_{0}),\quad z_{g}=\text{enc}(o_{g}).$
The MPC framework and CEM optimization procedure is detailed in Appendix A.4.1. Since our world model is differentiable, a possibly more efficient approach is to optimize this objective through gradient descent (GD), allowing the world model to directly guide the agent toward a specific goal. The details of GD are provided in Appendix A.4.2. However, we empirically observe that CEM outperforms GD in our experiments. We hypothesize this is due to our choice to not constrain the terrain smoothness of the world model during training, potentially leading to issues with the gradient. Full results for both planners can be found in Appendix A.4.3.
@@ 4 Experiments
Our experiments are designed to address the following key questions: 1) Can we effectively train DINO-WM using pre-collected offline datasets? 2) Once trained, can DINO-WM be used for visual planning? 3) To what extent does the quality of the world model depend on pre-trained visual representations? 4) Does DINO-WM generalize to new configurations, such as variations in spatial layouts and object arrangements? To answer these questions, we train and evaluate DINO-WM across five environment suites (full description in Appendix A.1) and compare it to a variety of state-of-the-art world models that model the world both in latent space and in raw pixel space.
@@ 4.1 Environments and Tasks
We consider five environment suites in our evaluations spanning simple navigation environments and manipulation environments with varying dynamics complexity. For all environments, the observation space is RGB images of size (224, 224).
Figure 3: We evaluate DINO-WM on 5 environment suites, from left to right: PointMaze, Push-T, Two Room, Rope Manipulation, and Granular Manipulation.
@@
    Point Maze: A simple 2D point maze navigation environment in the D4RL suite Fu et al. (2021). A point agent with 2-dimensional action space moves in a U-shape maze. The agent’s dynamics incorporate physical properties such as velocity, acceleration, and inertia, making the movement realistic. The objective of the task is to navigate the maze and reach arbitrary goal locations from arbitrary starting location.
@@
    Push-T: This manipulation environment was introduced in Chi et al. (2024) to study precise manipulation. The environment features a pusher agent interacting with a T-shaped block. The goal is to guide both the agent and the T-block from a randomly initialized state to a known feasible target configuration within 25 steps. The task requires both the agent and the T to match the target locations. Unlike previous setups, the fixed green T no longer represents the target position for the T-block but serves purely as a visual anchor for reference. Success requires precise understanding of the contact-rich dynamics between the agent and the object, making it a challenging test for visuomotor control and object manipulation. We also introduce a variant of this where we have multiple object shapes.
@@
    Wall: This custom 2D navigation environment featuring two rooms separated by a wall with a door opening. The task requires the agent to navigate from a randomized starting location in one room to a goal in the other room, which requires the agent to pass through the door. We introduce a variant of this environment where the positions of the wall and door are randomized to assess the model’s ability to generalize to novel configurations of familiar environment dynamics.
@@
    Rope Manipulation: This task is simulated with Nvidia Flex Zhang et al. (2024) and consists of an XArm interacting with a rope placed on a tabletop. The objective is to move the rope from an arbitrary start configuration to a specified goal configuration.
  5. e)
    Granular Manipulation: Granular manipulation uses the same setting as Rope manipulation and manipulates about a hundred particles to form desired shapes.
4.2 Baselines
We compare DINO-WM with the following state-of-the-art models commonly used for control:
  1. a)
    IRIS Micheli et al. (2023): IRIS employs a discrete autoencoder to translate visual inputs into tokens, and a GPT Transformer that predicts tokens of future observations. It combines these components to learn policies and value functions through imaginative procedures.
  2. b)
    DreamerV3 Hafner et al. (2024): DreamerV3 learns a world model to interpret visual inputs into categorical representation. It predicts future representations and rewards based on given action and trains an actor-critic policy from its imagined trajectories.
  3. c)
    TD-MPC2 Hansen et al. (2024) : TD-MPC2 learns a decoder-free world model in latent space and uses reward signals to optimize the latents. It serves as a strong baseline for reconstruction-free world modeling.
  4. d)
    AVDC Ko et al. (2023): AVDC leverages a diffusion model to generate an imagined video of task execution based on initial observation and a textual goal description. It then estimates optical flow between frames to capture object movements and generates robot arm commands.
Figure 4: Openloop rollout of world models trained with various pre-trained encoders on Push-T and Granular environment. For each trajectory, the model is given the first frame as well as sequence of actions. The world models performs openloop rollout with these actions, and the images are reconstructed by a pre-trained decoder. For each environment, the bottom row denotes the ground truth. DINO-WM (Ours) rollouts are bolded and are visually indistinguishable from the ground truth observations.
@@ 4.3 Optimizing Behaviors with DINO-WM
With a trained world model, we study if DINO-WM be used for zero-shot planning directly in the latent space.
For the PointMaze, Push-T, and Wall environments, we sample 50 initial and goal states to measure the success rate across all instances. Due to the environment stepping time for the Rope and Granular environments, we evaluate the Chamfer Distance (CD) on 10 instances for them. In the Granular environment, we sample a random configuration from the validation set, with the goal of pushing the materials into a square shape at a randomly selected location and scale.
Table 1: Planning results for offline world models on five control environments.
Model PointMaze PushT Wall Rope Granular SR $\uparrow$ SR $\uparrow$ SR $\uparrow$ CD $\downarrow$ CD $\downarrow$ IRIS 0.74 0.32 0.04 1.11 0.37 DreamerV3 1.00 0.04 1.00 2.49 1.05 TD-MPC2 0.00 0.00 0.00 2.52 1.21 Ours 0.98 0.90 0.96 0.41 0.26
As seen in Table 1, on simpler environments such as Wall and PointMaze, DINO-WM is on par with state-of-art world models like DreamerV3. However, DINO-WM significantly outperforms prior work at manipulation environments where rich contact information and object dynamics need to be accurately inferred for task completion. We notice that for TD-MPC2, the lack of reward signal makes it difficult to learn good latent representations, which subsequently results in poor performance. Visualizations of some planning results can be found in Figure 5.
@@ 4.4 Does pre-trained visual representations matter?
We use different pre-trained general-purpose encoders as the observation model of the world model, and evaluate their downstream planning performance. Specifically, we use the following encoders commonly used in robotics control and general perception: R3M Nair et al. (2022), ImageNet pretrained ResNet-18 Russakovsky et al. (2015); He et al. (2016) and DINO CLS Caron et al. (2021). Detailed descriptions of these encoders are in Appendix A.3.
@@
Model PointMaze PushT Wall Rope Granular SR $\uparrow$ SR $\uparrow$ SR $\uparrow$ CD $\downarrow$ CD $\downarrow$ R3M 0.94 0.42 0.34 1.13 0.95 ResNet 0.98 0.2 0.12 1.08 0.90 DINO CLS 0.96 0.44 0.58 0.84 0.79 DINOPatch (Ours) 0.98 0.90 0.96 0.41 0.26
@@ 4.5 Generalizing to Novel Environment Configurations
We would like to measure the generalization capability of our world models not just across different goals in an environment, but across different environments themselves. For this we construct three families of environments, where the world model will be deployed in an unseen environment for unseen goals. Our families of environments consist of WallRandom, PushObj, and GranularRandom with detailed descriptions in Appendix A.2. Visualizations of training and testing examples are shown in Figure 6.
Figure 6: Training and testing visualizations for WallRandom, PushObj and GranularRandom. Test setups are highlighted in blue boxes, showcasing unseen configurations for assessing the model’s generalization ability.
@@
From Table 3, we observe that DINO-WM demonstrates significantly better performance in the WallRandom environment, indicating that the world model has effectively learned the general concepts of walls and doors, even when they are positioned in locations unseen during training. In contrast, other methods struggle to accurately identify the door’s position and navigate through it. The PushObj task remains challenging for all methods, as the model was only trained on the four object shapes, which makes it difficult to infer physical parameters like the center of gravity and inertia precisely. In GranularRandom, the agent encounters fewer than half the particles present during training, resulting in out-of-distribution images compared to the training instances. Nevertheless, DINO-WM accurately encodes the scene and successfully gathers the particles into a designated square location with the lowest Chamfer Distance (CD) compared to the baselines, demonstrating better scene understanding. We hypothesize that this is due to DINO-WM’s observation model encoding the scene as patch features, making the variance in particle number still within the distribution for each image patch.
@@ 4.6 Qualitative comparisons with generative video models
Given the prominence of generative video models, it is reasonable to presume that they could readily serve as world models. To investigate the usefulness of DINO-WM over such video generative models, we compare it with imagined rollouts from AVDC Ko et al. (2023), a diffusion-based generative model. As seen in Figure 7, we find that the diffusion model trained on benchmarks produce future images that are mostly visually realistic, however they are not physically plausible as we can see that large changes can occur in a single timestep of prediction, and may have difficulties in reaching to the exact goal state. Potentially stronger generative models in the future could alleviate this issue.
We also compare DINO-WM with a variant of AVDC, where the diffusion model is trained to generate the next observation $o_{t+1}$ conditioned on the current observation $o_{t}$ and action $a_{t}$, rather than generating an entire sequence of observations at once conditioned on a text goal. As detailed in Appendix A.5, it can be seen that the action-conditioned diffusion model diverges from the ground truth observations over long-term predictions, making it insufficient for accurate task planning.
Figure 7: Comparison of plans generated by DINO-WM and AVDC, a diffusion-based generative model.
@@ 4.7 Decoding and Interpreting the Latents
Although DINO-WM operates in latent space and the observation model is not trained with pixel reconstruction objectives, training a decoder is still valuable for interpreting the model’s predictions. We evaluate the image quality of predicted futures across all models and find that our approach outperforms others, even those whose encoders are trained with environment-specific reconstruction objectives. We show openloop rollout visualizations in Figure 4. This demonstrates the robustness of DINO-WM despite the lack of explicit pixel-level supervision. We report two key metrics: Structural Similarity Index (SSIM) Wang et al. (2004) and Learned Perceptual Image Patch Similarity (LPIPS) Zhang et al. (2018) on the reconstruction of world models’ predicted future states. SSIM measures the perceived quality of images by evaluating structural information and luminance consistency between predicted and ground-truth images, with higher values indicating greater similarity. LPIPS, on the other hand, assesses perceptual similarity by comparing deep representations of images, with lower scores reflecting closer visual similarity.
LPIPS $\downarrow$ SSIM $\uparrow$
@@
Method PushT Wall Rope Granular PushT Wall Rope Granular
R3M 0.045 0.0083 0.023 0.08 0.956 0.994 0.982 0.917
ResNet 0.063 0.0024 0.025 0.08 0.950 0.996 0.980 0.915
DinoCLS 0.039 0.004 0.029 0.086 0.973 0.996 0.980 0.912
AVDC 0.046 0.030 0.060 0.106 0.959 0.983 0.979 0.909
Ours 0.007 0.0016 0.009 0.035 0.985 0.997 0.985 0.940
5 Conclusion, Limitations & Future Work
@@
In this work, we introduce DINO-WM, a simple yet effective technique for modeling visual dynamics in latent space without the need for pixel-space reconstruction. We have demonstrated that DINO-WM captures environmental dynamics and generalizes to unseen configurations, independent of task specifications, enabling visual reasoning at test time and generating zero-shot solutions for downstream tasks through planning. DINO-WM takes a step toward bridging the gap between task-agnostic world modeling and reasoning and control, offering promising prospects for generic world models in real-world applications. For limitations, DINO-WM still relies on the availability of ground truth actions from agents, which may not always be feasible when training with vast video data from the internet. Additionally, while we currently plan in action space for downstream task solving, an extension of this work could involve developing a hierarchical structure that integrates high-level planning with low-level control policies to enable solving more fine-grained control tasks.
Ethics Statement
This work explores creation of latent world models that can be used for better downstream planning. While we do not anticipate a potential for current misuse as this particular work, we can imagine future work that builds on this can lead to impact in robotics. Such potential applications to robotics open up a potential to misuse, which we acknowledge.
Reproducibility Statement
All code, models, and benchmarks produced from this project will be made open-source on our project website. We also provide thorough textual descriptions of all experimental procedures in the Appendix. Appendix A.1 describes our environments, data generation, and task definitions. In Appendix A.4, we outline all the planning optimization methods that we used in this paper. We provide further comparisons for DINO-WM and using generative models as world models in Appendix A.5, and additional planning visualizations with DINO-WM in Appendix A.7. Finally, Appendix A.6 provides the hyperparameters we used for training the world model for reproducing our experiment results in Section 4.1.

### v2

@@ 3.1.3 Decoder for Interpretability
To aid in visualization and interpretability, we use a stack of transposed convolution layers to decode the patch representations back to image pixels, similar as in (Razavi et al., 2019). Given a pre-collected dataset, we optimize the parameters $\theta$ of the decoder $q_{\theta}$ with a simple reconstruction loss defined as:
@@
The training of the decoder is entirely independent of the transition model training, offering several advantages: 1) The decoder does not affect the world model’s reasoning and planning capabilities for solving downstream tasks, and 2) There is no need to reconstruct raw pixel images during planning, thereby reducing computational costs. Nevertheless, the decoder remains valuable as it enhances the interpretability of the world model’s predictions. While backpropagating this decoder loss to the predictor is possible, we ablate this choice and find that it negatively impacts performance compared to omitting the decoder loss. Full details are provided in Appendix A.4.3.
@@ 3.2 Visual Planning with DINO-WM
To evaluate the quality of the world model, we perform trajectory optimization at test time and measure performance. While the planning methods themselves are fairly standard, they serve as means to emphasize the quality of the world models. For this purpose, our world model receives the current observation $o_{0}$ and a goal observation $o_{g}$, both represented as RGB images. We formulate planning as the process of searching for a sequence of actions that the agent would take to reach $o_{g}$. We employ model predictive control (MPC), which facilitates planning by considering the outcomes of future actions.
We utilize the cross-entropy method (CEM) to optimize the sequence of actions at each iteration. The planning cost is defined as the mean squared error (MSE) between the current latent state and the goal’s latent state, given by
$\mathcal{C}=\left\|\hat{z}_{T}-z_{g}\right\|^{2},\quad\text{where}\quad\begin{aligned} \hat{z}_{t}&=p(\hat{z}_{t-1},a_{t-1}),\\ \hat{z}_{0}&=\text{enc}(o_{0}),\\ z_{g}&=\text{enc}(o_{g}).\end{aligned}$
The MPC framework and CEM optimization procedure are detailed in Appendix A.5.1. Since our world model is differentiable, a possibly more efficient approach is to optimize this objective through gradient descent (GD), allowing the world model to directly guide the agent toward a specific goal. The details of GD are provided in Appendix A.5.2. However, we empirically observe that CEM outperforms GD in our experiments with full results in Appendix A.5.3. We hypothesize that incorporating regularizations during training and in the planning objectives could further improve performance, and leave this for future work.
@@ 4 Experiments
Our experiments are designed to address the following key questions: 1) Can we effectively train DINO-WM using precollected offline datasets? 2) Once trained, can DINO-WM be used for visual planning? 3) To what extent does the quality of the world model depend on pre-trained visual representations? 4) Does DINO-WM generalize to new configurations, such as variations in spatial layouts and object arrangements? We train and evaluate DINO-WM across six environment suites (full description in Appendix A.1) and compare it to a variety of state-of-the-art world models that predict in either latent space or raw pixel space.
@@ 4.1 Environments and Tasks
We evaluate six environment suites with varying dynamics complexity, some of which are drawn from standard robotics benchmarks, such as D4RL (Fu et al., 2021) and DeepMind Control Suite (Tassa et al., 2018), as shown in Figure 3. These environments include maze navigation (Maze, Wall), fine-grained control for tabletop pushing (PushT) and robotic arm control (Reach), and deformable object manipulation with an XArm (Rope, Granular).
In all environments, the task is to reach a randomly sampled goal state specified by a target observation, starting from arbitrary initial states. For PushT, target configurations are sampled to ensure feasibility within 25 steps. For Granular, targets require gathering all particles into a square with randomized locations and sizes. Observations in all environments are RGB images of size (224, 224). A full description of the environments is provided in Appendix A.1.
Figure 3: We evaluate DINO-WM on six environment suites, from left to right, top to bottom: Maze, Reach, Wall, Push-T, Rope Manipulation, and Granular Manipulation.
4.2 Baselines
We compare DINO-WM with the following state-of-the-art models commonly used for control. For IRIS, DreamerV3, and TD-MPC2, we train the models with our offline datasets without any reward or task information, and perform MPC on the learned world model for solving downstream tasks.
@@ 4.2 Baselines
    IRIS (Micheli et al., 2023): IRIS encodes visual inputs into tokens via a discrete autoencoder and predicts future tokens using a GPT Transformer, enabling policy and value learning through imagination.
@@
    DreamerV3 (Hafner et al., 2024): DreamerV3 encodes visual inputs into categorical representations, predicts future states and rewards, and trains an actor-critic policy from imagined trajectories.
@@
    TD-MPC2 (Hansen et al., 2024) : TD-MPC2 learns a decoder-free world model in latent space and uses reward signals to optimize the latents.
@@
    AVDC (Ko et al., 2023): AVDC uses a diffusion model to generate task execution videos from an initial observation and textual goal. We provide qualitative evaluations and MPC planning results for an action-conditioned variant in Appendix A.6.
Figure 4: Open-loop rollouts of world models on Push-T and Granular. Given the first frame and action sequence, each model predicts future frames, reconstructed by its decoder. For each environment, the bottom row denotes the ground truth. DINO-WM (Ours) rollouts are bolded and are visually indistinguishable from the ground truth observations.
@@ 4.3 Optimizing Behaviors with DINO-WM
With a trained world model, we study if DINO-WM can be used for zero-shot planning directly in the latent space.
For Maze, Reach, PushT, and Wall environments, we sample 50 initial and goal states and measure the success rate across all instances. Due to the environment stepping time for the Rope and Granular environments, we evaluate the Chamfer Distance (CD) on 10 instances for them. In Granular, we sample a random configuration from the validation set, with the goal of pushing the materials into a square shape at a randomly selected location and scale.
Table 1: Planning results for offline world models on six control environments.
Model Maze Wall Reach PushT Rope Granular SR $\uparrow$ SR $\uparrow$ SR $\uparrow$ SR $\uparrow$ CD $\downarrow$ CD $\downarrow$ IRIS 0.74 0.04 0.18 0.32 1.11 0.37 DreamerV3 1.00 1.00 0.64 0.30 2.49 1.05 TD-MPC2 0.00 0.00 0.00 0.00 2.52 1.21 Ours 0.98 0.96 0.92 0.90 0.41 0.26
As seen in Table 1, on simpler environments such as Wall and PointMaze, DINO-WM is on par with state-of-the-art world models like DreamerV3. However, DINO-WM significantly outperforms prior work at manipulation environments where rich contact information and object dynamics need to be accurately inferred for task completion. We notice that for TD-MPC2, the lack of reward signal makes it difficult to learn good latent representations, which subsequently results in poor performance. Visualizations of planning on all environments can be found in Appendix A.10.
Does DINO-WM learn better environment dynamics as more data become available? We conduct a set of ablation experiments in Appendix A.4.1, showing that the planning performance scales positively with the amount of training data. We also present the full inference and planning times for DINO-WM in Appendix A.8, showing significant speedup over traditional simulation, particularly in the computationally intensive deformable environments.
@@ 4.4 Does pre-trained visual representations matter?
We use different pre-trained general-purpose encoders as the observation model of the world model, and evaluate their downstream planning performance. Specifically, we use the following encoders commonly used in robotics control and general perception: R3M (Nair et al., 2022), ImageNet pretrained ResNet-18 (Russakovsky et al., 2015; He et al., 2016) and DINO CLS (Caron et al., 2021). Detailed descriptions of these encoders are in Appendix A.3.
@@
Model Maze Wall Reach PushT Rope Granular SR $\uparrow$ SR $\uparrow$ SR $\uparrow$ SR $\uparrow$ CD $\downarrow$ CD $\downarrow$ R3M 0.94 0.34 0.40 0.42 1.13 0.95 ResNet 0.98 0.12 0.06 0.20 1.08 0.90 DINO CLS 0.96 0.58 0.60 0.44 0.84 0.79 DINOPatch (Ours) 0.98 0.96 0.92 0.90 0.41 0.26
@@ 4.5 Generalizing to Novel Environment Configurations
We evaluate the generalization of our world models not only across different goals but also across various environment configurations. We construct three environment families—WallRandom, PushObj, and GranularRandom—where the model is tested on unseen configurations with random goals. Detailed descriptions of the environments can be found in Appendix A.2.
Figure 5: Training and testing visualizations for WallRandom, PushObj and GranularRandom. Test setups are highlighted in blue boxes.
@@
From Table 5, we observe that DINO-WM demonstrates significantly better performance in WallRandom, indicating that model has effectively learned the general concepts of walls and doors, even when they are positioned in locations unseen during training. In contrast, other methods struggle to accurately identify the door’s position and navigate through it. The PushObj task remains challenging for all methods, as the model was only trained on the four object shapes, which makes it difficult to precisely infer relevant physical parameters. In GranularRandom, the agent encounters fewer than half the particles present during training, resulting in out-of-distribution images compared to the training instances. Nevertheless, DINO-WM accurately encodes the scene and successfully gathers the particles into a designated square location with the lowest Chamfer Distance (CD) compared to the baselines, demonstrating better scene understanding. We hypothesize that this is due to DINO-WM’s observation model encoding the scene as patch features, making the variance in particle number still within the distribution for each image patch.
@@ 4.6 Qualitative comparisons with generative video models
Given the prominence of generative video models, it’s natural to assume they could serve as world models. We compare DINO-WM with AVDC (Ko et al., 2023), a diffusion-based generative model. As shown in Figure 6, while AVDC can generate visually realistic future images, these images lack physical plausibility. Large, unrealistic changes can occur within a single timestep, and the model struggles to reach the exact goal state. Future advancements in generative models may help address these issues.
We further compare DINO-WM with a variant of AVDC, where the diffusion model is trained to generate the next observation $o_{t+1}$ conditioned on the current observation $o_{t}$ and action $a_{t}$. As detailed in Appendix A.6, the action-conditioned diffusion model diverges from the ground truth observations over long-term predictions, making it insufficient for accurate task planning.
Figure 6: Plans generated by DINO-WM and AVDC.
@@ 4.7 Decoding and Interpreting the Latents
Although DINO-WM operates in latent space and the observation model is not trained with pixel reconstruction objectives, training a decoder aids in interpreting predictions. We evaluate the image quality of predicted futures across all models and find that our approach outperforms others, even those whose encoders are trained with environment-specific reconstruction objectives. Open-loop rollouts in Figure 4 demonstrate DINO-WM’s robustness despite the lack of explicit pixel supervision. We report the Learned Perceptual Image Patch Similarity (LPIPS) (Zhang et al., 2018) on the world models’ predicted future frames, which assesses perceptual similarity by comparing deep representations of images, with lower scores reflecting closer visual similarity. Additional results, including Structural Similarity Index (SSIM) (Wang et al., 2004), are provided in Appendix A.7.
Table 4: Comparison of world models on LPIPS metrics.
@@
Method PushT Wall Rope Granular R3M 0.045 0.008 0.023 0.080 ResNet 0.063 0.002 0.025 0.080 DINO CLS 0.039 0.004 0.029 0.086 AVDC 0.046 0.030 0.060 0.106 Ours 0.007 0.0016 0.009 0.035
5 Conclusion
@@
We introduce DINO-WM, a simple yet effective technique for modeling visual dynamics in latent space without the need for pixel-space reconstruction. We have demonstrated that DINO-WM captures environmental dynamics and generalizes to unseen configurations, independent of task specifications, enabling visual reasoning at test time and generating zero-shot solutions for downstream tasks through planning. DINO-WM takes a step toward bridging the gap between task-agnostic world modeling and reasoning and control, offering promising prospects for generic world models in real-world applications.
Limitations and Future Work: First, DINO-WM assumes having access to offline datasets with sufficient state-action coverage, which can be challenging to obtain for highly complex environments. This can potentially be addressed by combining DINO-WM with exploration strategies and updating the model as new experiences are available. Second, DINO-WM still relies on the availability of ground truth actions from agents, which may not always be feasible when training with vast video data from the internet. Lastly, while we currently plan in action space for downstream task solving, an extension of this work could involve developing a hierarchical structure that integrates high-level planning with low-level control policies to enable solving more fine-grained control tasks.
Impact Statement
This paper presents work whose goal is to facilitate the learning and applications of task-agnostic world models. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
Acknowledgements
We would like to thank Ademi Adeniji, Alfredo Canziani, Amir Bar, Kevin Zhang, Mido Assran, Vlad Sobal, Zichen Jeff Cui for their valuable discussion and feedback. This work was supported by grants from Honda, Hyundai, NSF award 2339096 and ONR awards N00014-21-1-2758 and N00014-22-1-2773. LP is supported by the Packard Fellowship.